Two point charges 16nC and 8nC are situated at the corners B and C of an equilateral triangle of side 0.03m. Find the magnitude and direction of the resultant electric field at the vertex A of the triangle.

Grade:12

1 Answers

Ash
5 years ago
The electric field due to the charge 16nc,
E1=\frac{1}{4\pi \varepsilon }[\frac{q1}{r^{2}}]
E1=16x10^4 away from the charge
Similarly,
The electric field due to the point charge 8nc,
E2=8x10^4 away from the charge
the resultant field
E=\sqrt{E1^{2}+E2^{2}+2E1E2cos\Theta }       (angle=60 degrees)
E=21.1x10^4 N/C
Direction: Along and away from the perpendicular bisector of BC